Problem 5A-8 High-Low Method; Predicting Cost [LO5-10]

Nova Company’s total overhead cost at various levels of activity are presented below:

Month Machine-
Hours
Total
Overhead
Cost
April 52,000 $ 207,440
May 42,000 $ 181,740
June 62,000 $ 233,140
July 72,000 $ 258,840

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 42,000 machine-hour level of activity is:

Utilities (variable) $ 63,000
Supervisory salaries (fixed) 56,000
Maintenance (mixed) 62,740
Total overhead cost $ 181,740

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Required:

1. Estimate how much of the $258,840 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $258,840 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.)

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX.

3. Express the company’s total overhead cost in the form Y = a + bX.

4. What total overhead cost would you expect to be incurred at an activity level of 47,000 machine-hours?
Answer #1
1) Utilities (63000/42000)*72000 108000
Supervisory Salaries 56,000
Maintenance (balancing figure) 94,840
Total overhead cost 258,840
Maintenance cost in july 94,840
2) Estimating formula for maintenance cot
MH's cost
high 72,000 94,840
low 42,000 62,740
Difference 30,000 32,100
Variable cost per unit = 32100/30000
1.07
Fixed cost = 94840-72000*1.07
17800
Cost formula
Y= 17,800 + 1.07 X
3) Companys total overhead cost
Variable Fixed
Utilities 1.3
Supervisory salaries 56,000
Maintenance cost 1.07 17,800
total 2.37 73800
cost formula
Y= 73,800 + 2.37 X
4) total overhead cost
Y= 73,800 + 2.37 X
so
fixed cost 73,800
Variable cost (47000*2.37) 99540
total overhead cost 173,340 answer